OpenStudy (anonymous):

If it was 50°F outside, what was the temperature in degrees Celsius?

OpenStudy (rmrjr22):

10 degrees celcius

OpenStudy (rmrjr22):

(°F - 32) x 5/9 = °C

OpenStudy (anonymous):

use C=5(F-32) /9 =5(50-32)/9=5*18/9 =10

OpenStudy (anonymous):

thank you ;)

OpenStudy (rmrjr22):

Best response? :)